two-thirds of Helen's age is half of Dale's age. Dale is 10 years older than Helen. How old is Helen?
Answer by jorel1380(3719)

d=h+10;d-10=h
2h/3=d/2
2h/3=h+10/2
4h=3h+30
h=30
d=40
Helen is 30. Dale is 40..